OpenStudy (anonymous):

Solve for zeros: y= -2 (x+3)(x-7)

OpenStudy (anonymous):

anyone :(

OpenStudy (callisto):

I hope I didn't misunderstand the question.. Put y=0 into the equation 0 = -2 (x+3) (x-7) Divide both sides by -2 0 = (x+3)(x-7) x+3 = 0 or x-7 = 0 <- solve these two equations

OpenStudy (anonymous):

THANKYOU !

OpenStudy (callisto):

Welcome :)